Cumene is a compound containing only carbon and hydrogen. It is 89.94% C by mass and has a molar mass of 120.2 g/mol. Determine the empirical and molecular formulas of the compound.
1
Expert's answer
2019-09-25T05:45:56-0400


m(C)= 120.2 g x 89.94% = 108.1 g 

m(H)=120.2 g -108.1 g = 12.09 g  

n(C)=108.1 / 12.01 = 9.000 mol 

n(H)=12.09 / 1.008 = 11.99 mol 

“C” 9.000 / 9.000 = 1 

“H” 11.99 / 9.000 = 1.332 

The empirical formula of the substance is (C3H4)n 

M(C3H4)=40.06g/mol

M(Cumene(C9H12))=120.2 g/mol

120.2/40.06=~3

The molecular formula is then 3 x C3H4 or C9H12.